The weight of copper anode and zinc anode decreases . Explain

Grade:11

1 Answers

Usman Iqbal
6 years ago
At anode oxidation takes place , so zinc and copper anode undergoes oxidation as follows : Zn = Zn+2 + 2electronsCu = Cu+2 + 2 electrons.Thus Zinc and Copper anode keeps getting thinner as they oxideses to form dissolved Zn+2 and Cu+2 ions respectively.